OpenStudy (anonymous):

Determine if the situation is best modeled by a linear model or an exponential model, and then write the equation that represents the following situation: Allison’s pay if she earns $5 for each hour she babysits. Answer linear; y = 5x exponential; y = 5x linear; y = 5^x exponential; y = 5^x

OpenStudy (rsadhvika):

look at the options we can strike off few options straight away

OpenStudy (rsadhvika):

exponential; y = 5x REJECT cuz, y=5x is linear

OpenStudy (rsadhvika):

linear; y = 5^x REJECT cuz, y=5^x is exponential

OpenStudy (rsadhvika):

so, answer must be either A or D

OpenStudy (anonymous):

is it A

OpenStudy (rsadhvika):

Correct it is A, cuz 5 +5+5... it is

OpenStudy (anonymous):

ya